[The following information applies to the questions displayed below.]

MoveIt Corporation is the world’s leading express-distribution company. In addition to its 643 aircraft, the company has more than 57,000 ground vehicles that pick up and deliver packages. Assume that MoveIt sold a delivery truck for $21,000. MoveIt had originally purchased the truck for $35,000 and had recorded depreciation for three years.

Required:

  1. Calculate the amount of gain or loss on disposal, assuming that Accumulated Depreciation was (a) $14,000, (b) $11,000, and (c) $19,000. (Select "None" if there is no Gain or Loss.)
Answer #1

Calculate gain or loss

Situation Cost Accumulated depreciation Book value Gain (loss)
1 35000 14000 21000 21000-21000 = 0
2 35000 11000 24000 21000-24000 = -3000 (loss)
3 35000 19000 16000 21000-16000 = 5000 Gain